40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4901     Entry time: Tue Jun 28 16:52:37 2011
Author: Sonali 
Type: Update 
Category: Green Locking 
Subject: Routing of fibre to PSL complete. 

1. Suresh and I completed the alignment of the fibre and the three mirrors on the ETMY table.

2. We managed to get an output beam power of around 60% using the Ophir(Orion/PD) power meter to finetune the alignment. The power of the input beam is 74.4 mW and of the output beam is 38.5 mW.

3. The coupler on the output side of the fibre which had been put there to help in the alignment has been removed.

4. The picture of the ETMY layout as of now has been attached.

5. The labels A stands for the mirror used to turn the beam direction and B and C stand for the three mirrors used in the alignment of the beam into the coupler,D.(attachment 3).

6. The fibre we used is 50m in length which was barely sufficient to reach the PSL table.

7. So, the fibre has been routed to the PSL table using the fibre tray running below the Y-arm tube as this was the shortest route possible(even though it is a rather acccident prone zone).

8. The fibre has been tied down at regular intervals so that it does not get snagged and pulled up inadvertently.

9. We will start with the preparation of the layout of the PSL table to superpose the two beams on Monday.

Attachment 1: coupled_fibre.jpg  3.077 MB  | Hide | Hide all
coupled_fibre.jpg
Attachment 2: the_fibre_route.jpg  2.682 MB  | Hide | Hide all
the_fibre_route.jpg
Attachment 3: ETMY_aftr_fibre_coupling2.png  5.510 MB  Uploaded Thu Jun 30 00:29:53 2011  | Hide | Hide all
ETMY_aftr_fibre_coupling2.png
ELOG V3.1.3-